OpenStudy (anonymous):

Solve for x the equation x^4 + ax^3 - ab^2 -a^2b^2 = 0 here a,b, are considered KNOWN parameters

OpenStudy (anonymous):

Answer in 3 minutes

OpenStudy (anonymous):

Hint: factor in two parentheses

OpenStudy (anonymous):

U posted the wrong question

OpenStudy (anonymous):

Yeah. You missed an x somewhere.

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

\[x^3(x+a)-ab^2(1+a)=0\] \[x=1,x^3=ab^2\]

OpenStudy (anonymous):

If u will factor this x^4 + ax^3 - ab^2 -a^2b^2 = 0 in 3years I MUST SAY U R GREAT

OpenStudy (anonymous):

@Jonask what r u doing?

OpenStudy (shubhamsrg):

just realised,,its easier than it looks,, direct ans is what @Jonask said.. which leads to what @sauravshakya said there and ultimately what @mukushla said,,hmm.. the range of my apologies inc,,ques should have ended 50-60 posts ago! hmm..

OpenStudy (anonymous):

@Jonask Great - simple, and final.

OpenStudy (anonymous):

Can u PLZ explain me

OpenStudy (shubhamsrg):

can you pin point there who said you did wrong and exactly what he said ? o.O

OpenStudy (anonymous):

Oh got it

OpenStudy (anonymous):

Explain what - the equation is in front of you SIR.

OpenStudy (anonymous):

Howw??? Put x = 1. => (1+a) = ab^2(1+a) ... How is that true???

OpenStudy (anonymous):

@Mikael , "SIR"? Explain

OpenStudy (anonymous):

Check @Jonask

OpenStudy (shubhamsrg):

not 1,,he must have meant -a .. @Mikael posted the incorrect eqn..

OpenStudy (shubhamsrg):

abe kitna bada gadha hai tu !! attitude problem!! hmm.. :(

OpenStudy (anonymous):

^^ Agreed, man.

hartnn (hartnn):

and why x can't be negative ? is it length of something? what ?

OpenStudy (anonymous):

all is correct vonjuch. dolbo-by

OpenStudy (shubhamsrg):

it corresponds to this silly thing : http://openstudy.com/study#/updates/5051f753e4b02b4447c15384 @hartnn

OpenStudy (anonymous):

This is from some geometric beginning @hartnn sorry

OpenStudy (shubhamsrg):

forgive him @hartnn !! :D

OpenStudy (anonymous):

CAN SOMEONE EXPALIN ME WHAT @Jonask did

OpenStudy (anonymous):

HE FACTORIZED (what was said)

OpenStudy (anonymous):

What he did is wrong. The question is. The link Mikael gave has the correct question, and the correct, solution.

hartnn (hartnn):

ok,ok.

OpenStudy (shubhamsrg):

@Jonask probably assumed it to be x^4 + ax^3 - Xab^2 -a^2b^2 = 0 now its easily done..

OpenStudy (anonymous):

How is @Jonask correct @Mikael

OpenStudy (anonymous):

He solved the INTENDED equation @sauravshakya

OpenStudy (shubhamsrg):

kya bolta hai,khud hi bol ke khud hi samjh ke baith jata hai !! am not gonna translate this one! :P

OpenStudy (anonymous):

factorize and according to equality,\[x^3(x+a)=ab^2(1+a)\] for the equation to be true\[x^3=ab^2,x=1\]

OpenStudy (shubhamsrg):

ohh,,since you did that,,,i'd its wrong..

OpenStudy (shubhamsrg):

i thought you by default assumed an x there..which made it correct with reference to http://openstudy.com/study#/updates/5051f753e4b02b4447c15384 @hartnn

OpenStudy (shubhamsrg):

ohh sorry,,@hartnn was not intended in previous post! :P

OpenStudy (shubhamsrg):

nevermind,,i guess you should close the case now @Mikael

OpenStudy (anonymous):

@siddhantsharan is correct in two remarks

OpenStudy (anonymous):

Mudach'jo chernogzopoe otsasyvaet dvupolo